The activity of a sample of the cesium isotope 137Cs is 150mCi . When delivered to a hospital 14 hours later, its activity is 95 mCi.

What is the isotope's half life?

If the minimum usable activity is 10 {\rm mCi}, how long after delivery at the hospital is the sample usable?
Answer #1

ln[A]?ln[A]o=?kt

ln[95]-ln[137]=-kx14hr

k=0.02615hr-1

t1/2=0.693/K=0.693/0.02615hr-1 =26.5 hours

ln[A]?ln[A]o=?kt

ln[10]?ln[150]o=-0.02615hr-1 t

t=103.6Hours will take for the sample to have activity 10 mCi

sample usuable after delivery=103.6-14=89.6Hours